Use the equation 12x-2y=16 to answer the following:
A)Write the equation in slope-intercept form.
B)Identify the slope of the equation.
C)Identify the y-intercept.

Answer by Alan3354(69443)

Use the equation 12x-2y=16 to answer the following:
A)Write the equation in slope-intercept form.
That's y = mx + b where m = slope, b = y-intercept
----------------------
That means solve for y
12x-2y=16
6x - y = 8
y = 6x - 8
--------------------
B)Identify the slope of the equation
m = 6
-------------
C)Identify the y-intercept
b = -8